an apple and an orange cost #250, the apple costs #200 more than the orange, how much does the orange cost?
Answer by ikleyn(52798)

.

    Orange + Apple   = 250


      x    + (x+200) = 250


        2x           = 250 - 200 = 50

         
         x                       = 50/2 = 25.


ANSWER.  1 orange costs #25;  one apple costs $225.

Solved.